OpenStudy (anonymous):

Integrals, integrals, integrals... Use Substitution formula to evaluate. see question in comment box ! :)

OpenStudy (anonymous):

\[\int\limits_{0}^{1} \frac{ 4x dx}{ \sqrt{4 +2x ^{2}} }\]

OpenStudy (anonymous):

what is the substitution formula ?

OpenStudy (sidsiddhartha):

substitute that

OpenStudy (anonymous):

I somehow doubt this is trig sub, sid.

OpenStudy (anonymous):

Well, I just assume this is u-substitution, just like before. So what do you think would be the appropriate choise for u?

OpenStudy (anonymous):

Wow, how the hell did I spell that choice as choise?! O_o

OpenStudy (anonymous):

you read my mind! okay u = \[\sqrt{4 + 2x ^{2}}\]

OpenStudy (anonymous):

Remember, we rarely want to choose a u-substitution that would be a big chain rule. That choice of u will create extra expressions of x that we dont want. Think of it this way, you usually always want u to be the inside of something. Inside of a power, like the (4+cosx)^6 problem from last time, we wanted u = 4+cosx, not the whole thing, inside of a square root, etc.

OpenStudy (anonymous):

du = 

OpenStudy (anonymous):

ok ok!!

OpenStudy (anonymous):

so u = 4 +2x^2

OpenStudy (anonymous):

Bingo.

OpenStudy (anonymous):

du= 4x

OpenStudy (anonymous):

du = 4xdx to be more specific :)

OpenStudy (anonymous):

i was getting there :)

OpenStudy (anonymous):

okay now what ? how many times before I get integrals eeek!!

OpenStudy (anonymous):

i know du/u

OpenStudy (anonymous):

Well, we have a u and we have an expression for dx u = 4 + 2x^2 du = 4xdx. That works out perfectly